# ]Meta Modern Recommendation System (MRS) RecoWorld: Building Simulated Environments for Agentic Recommender Systems

### v1

@@ Abstract
Fei Liu at , Xiangjun Fan at
1 Introduction
@@
Example Use Cases of RecoWorld | |
Case 1 | Evaluating RecSys’ Instruction Following Ability. Instruction-following capabilities are essential for agentic RecSys, much like they are for LLMs. E.g., a user might tell the recommender system their preferences through natural language or voice, e.g., “There are too many recommendations about hairstyling; I wanna see something different but related,” or “I’d like to watch more content that people in the San Francisco Bay Area are watching right now.” If the agentic RecSys adjusts its recommendations based on these instructions, we need a simulated environment to evaluate how well it performed. One approach is to measure the system’s success in instruction-following by using simulated user feedback as a proxy. This approach is important because agentic recommenders will need to handle a wide range of user instructions. Our RecoWorld provides a promising way to estimate system performance without relying on ground-truth annotated data. |
Case 2 | Enabling Creators to Experiment with Publishing Strategies. A simulated environment helps creators explore different content strategies and reach their goals, e.g., growing followers, boosting views, or getting more likes. It lets them try bold, different approaches without risking real user experience. For example, they can test new topics and styles to attract fresh audiences, find the right posting frequency to keep followers interested without annoying them, or see how controversial opinions might be received. By mimicking user feedback, this environment lets creators test the waters safely. Simulated user profiles can be based on their current audience, a general population sample, or even specific target groups. |
Case 3 | Supporting New and Marginal Users in Exploring Interests. Algorithms such as the multi-armed bandit are frequently used in RecSys to find the right balance between exploring new user interests and exploiting known preferences. I.e., an agent chooses from a fixed set of actions (interests) to maximize cumulative reward over time. A challenge here is to explore without compromising user experience. RecoWorld addresses this by collecting simulated user engagement metrics, such as watch time and clicks, which serve as “pseudo-rewards” to evaluate the relevance of each interest. Our simulator can seamlessly integrate with contextual bandit by simulating groups of users with similar features, such as age, location, and past engagements, to provide collective exploration feedback. |
Case 4 | Building a Community Leaderboard for Agentic RecSys. Many enterprise recommender systems are developed in-house, making it difficult to replicate their results. The challenge is even greater with Agentic RecSys, which can take on various forms. They share an “agentic” nature, e.g., an agent (or multiple agents) interacts with an environment; the agent takes actions and receives observations and rewards in return. However, each task may have a different setup, e.g., variations in environmental states and reward signals. RecoWorld provides a shared platform with a community leaderboard, allowing different systems to be compared fairly. We invite practitioners to contribute their agentic tasks to RecoWorld using a common API. They can receive feedback signals from our simulated users and make their tasks available for the community to participate in. RecoWorld thus accelerates RL development, organizes various recommendation tasks from simple to complex for curriculum learning, and even fosters an agent-agent collaborative ecosystem. |
@@
We introduce RecoWorld, a blueprint for building simulated environments that propel the development of agentic recommender systems (Huang et al., 2025a). As these systems gain capabilities such as perception, reasoning, planning, memory, tool use, and autonomy, having environments for rapid system testing and refinement becomes essential before deploying them widely. We do not propose new agentic recommenders here; instead, our focus is on creating simulated environments that replicate user feedback. RecoWorld harnesses LLMs’ strong reasoning, long-context understanding, and multimodal abilities to simulate user responses. It focuses on optimizing for long-term value through simulated interactions. The approach also emphasizes collective impacts, e.g., a post is distributed to a community of interconnected users rather than isolated individuals. The design of user simulation and reward signals in RecoWorld offers promising potential for enhancing existing recommender systems in modeling both short- and long-term user interests, as well as for generative recommenders. (Xia et al., 2025; Wang et al., 2025b).
@@
Our user simulator realistically models sequential user behavior. When a user is about to disengage, they can issue instructions, such as “show me more interesting content”, instead of simply dropping out. The agentic RecSys responds to these instructions by creating new item lists, aiming to re-engage the user. This cycle continues until the user chooses to exit without further instructions.
The simulator’s effectiveness can be conveniently evaluated using existing RecSys datasets and validated by human annotators. A successful simulator triggers accurate instructions that guide the recommender, while an effective recommender responds with item lists that enhance user satisfaction. Together, this leads to improved recommendation quality.
To our knowledge, RecoWorld is the first environment to enable instruction-following recommenders in a Gym-like RL framework, marking a significant step toward more interactive recommender systems.
2 Agentic Environments for RecSys
Agentic is emerging as a defining characteristic of next-generation foundation models. An agentic recommender system acts as an autonomous agent that actively learns from user interactions. It follows user instructions, acquires new skills to make proactive recommendations, and adapts its behavior based on these experiences (Silver and Sutton, 2025). This marks a shift from traditional recommender systems that primarily make passive suggestions (Huang et al., 2025a; Shang et al., 2025; Maragheh and Deldjoo, 2025). With new capabilities to reason and interact with users, there comes the urgent need to develop new environments for evaluation. Simulated environments enable rapid system iteration, enhance data curation, and are increasingly leveraged for agent training. For example, Claude-4 uses a simulated environment to check how the model responds to risky system prompts, such as being asked to act as a dark web shopping assistant (Anthropic, 2025). Kimi K2 introduces an agentic data synthesis pipeline that creates tool-use demonstrations via simulated and real environments, generating agentic interactions at scale (Kimi Team et al., 2025).
@@ 2 Agentic Environments for RecSys
Multi-Turn Interactions Within a Session. A simulated user interacts with an agentic RecSys over multiple turns (see Figure 1). The user receives a list of recommended items and can take actions such as clicking or skipping. In the first turn, the user clicks on two items (green), skips two items (yellow), and then chooses to leave the session after the next item (red). Before leaving, the user reflects on their experience, identifies reasons for dissatisfaction, and gives a short instruction to the RecSys as feedback. If the RecSys cannot process user instructions, it simply updates recommendations based on recent interactions. If it can, it refreshes the recommendations to better address the user’s feedback. The user reviews the new list, continues interacting, and may repeat the process of providing feedback if considering leaving. This cycle continues until the user decides to end the session without further feedback. The sequence of recommended items and the user’s actions form the interaction trajectory for the session.
Consider an example instruction: User: “Show me more interesting content.” System: “Absolutely! Here are videos tailored to your interests and engagement history. Would you like to specify what you find interesting, such as trending topics, educational content, or entertainment? You can also tell me if you want more from specific creators or communities.” The agent perceives the user’s explicit instruction and reasons over their engagement history, demographics, and behavior to create a configuration for the recommender system to execute. The system returns an updated recommendation list along with strategies to guide the user. The user simulator then generates feedback signals indicating whether the Agentic RecSys met the user’s instructions and may provide the next instruction.
@@
RecoWorld is designed to be highly flexible. (a) When simulated users are replaced by human annotators, the agentic RecSys interacts with actual users and receives authentic feedback signals, effectively functioning as an online experiment. (b) When Agentic RecSys is replaced with a traditional RecSys, the system ignores user instructions and generates recommendations only from the user’s past engagement history, without adapting to additional feedback. (c) When instructions are disabled, the user simulator defaults to an evaluator that scores a list of items, without giving instructions to improve. Since no improvement is requested, there will be no follow-up iterations; the session will end after a single turn.
Explicit Versus Implicit Instructions Importantly, user preferences are not conveyed solely through explicit instructions; they are also conveyed implicitly through behavior, whether users realize it or not. E.g., if a user frequently finishes watching videos that are >5 minutes, it indicates a receptiveness to long-form content, unlike those who prefer quick 10- or 20-second clips. If a user continues a session without interruption, it likely means they are satisfied with the recommendations. Conversely, if they request “something more interesting,” or if a user does not return for a while, it likely indicates that the previous recommendations failed to engage them. Implicit instructions can be inferred by monitoring user state. Users with similar behavioral patterns often share underlying implicit signals. Collaborative filtering can exploit these shared patterns to identify implicit instructions that reflect the collective interests of these user groups. Both implicit and explicit signals are crucial for agentic recommenders.
@@ 3 User Simulation
A session begins when the user opens the app and ends when they exit. Each recommendation presents a list of items, selected from a candidate set and shown in order. The agent and simulated user may interact over multiple turns within a session, e.g., the user requests more interesting content, and the agent updates the list. A reward signal is generated after each list is shown. Crucially, we aim to optimize for long-term user retention as the reward signal, i.e., maximizing session duration and minimizing the interval between sessions, which relates to Daily Active Users (DAU). Specifically, the simulated user goes through each recommendation one by one. For each item, the user takes three steps: 1. Think it through: What is your reasoning behind how you’ll respond? 2. Take action: Write down what you’ll actually do. 3. Update your mindset: How does this affect your current thinking? Figure 3 shows a summary table containing initial simulator results, including simulated user actions and reasoning for each item.
3.1 Multimodal User Engagment Modeling
@@
Text-based modeling. One approach is to represent all user and item data in textual form. User demographics can be specified in the system prompt, while each item’s metadata, such as publish time, genre, creator, content type, interest cluster, and a brief summary, together with the user’s interaction history, are provided as textual context. This approach offers flexibility in determining which information to include in lifelong user behaviors, supports prompt compression for improved efficiency, and leverages the reasoning ability of LLMs to analyze user behavior and replicate user responses. Nonetheless, relying solely on textual descriptions may overlook certain nuances, including the mood of a video, sarcasm in content, or musical characteristics, which can limit the fidelity of preference modeling.
Multimodal modeling. The second alternative is to use a multimodal LLM (MLLM) as the backbone, such as Qwen2.5’s Omni model (Xu et al., 2025b) or vision-language models (VLM), Gemini-2.5-Pro. In this setting, items are directly ingested by the model rather than represented through textual descriptions. Images are tokenized into patches, while short-form videos with their audio tracks can be jointly processed through frameworks that synchronize visual and acoustic signals. Demographic attributes and textual metadata can also be incorporated as auxiliary inputs. The model output may take multiple modalities, though we restrict it to text for predicting user behavior. Compared with text-only LLMs, multimodal models provide richer representational capacity for multimodal content, enabling more faithful input understanding. Nevertheless, their reasoning ability, typically weaker, and the substantially larger token space of audio–visual inputs imposes significant context and efficiency challenges, making prompt compression particularly difficult.
@@
In addition to effective multimodal user engagement modeling, it is essential to achieve robust lifelong user behavior modeling. Users continuously interact with diverse items in recommender systems, resulting in ever-growing interaction histories. Over time, these histories can become effectively unbounded, making it computationally infeasible and costly to store and process all past behaviors. Precisely, we highlight two crucial principles built upon multimodal modeling: 1. Dynamic memory modeling: Users may have infinite historical interactions. However, not all behaviors contribute equally to reflecting the user’s current intention, and many may introduce noise. It is therefore crucial to design mechanisms that can dynamically and intelligently store, filter, and retrieve user historical behaviors. 2. Evolving preference modeling: User interests are not static but continuously evolve with changing contexts and the information they are exposed to through recommendations. Therefore, a user simulator needs to effectively capture temporal dynamics, long-term dependencies, and potential shifts in user interests, ensuring that the generated behaviors and the resulting recommendations remain both personalized and timely throughout the interaction lifecycle. In our proposed method, this can be achieved via mindset updates (refer to step-by-step feedback example in Figure 3).
3.2 Candidate Items
@@
We use real-world content whenever possible (e.g., engagement history and recommended items) to ensure realistic user behavior modeling. Using both simulated content and simulated users can be challenging, as combining these approximations may affect the reliability of behavioral insights. Specifically, each item is presented by its content (e.g., short-form video) along with its metadata and text description. The metadata includes details such as the timestamp, creator information, engagement metrics (e.g., likes, shares, comments), and content category. A user may be recommended various types of content, such as: (a) short-form videos, (b) text posts, (c) mixed media (posts, videos, and images), and (d) friend suggestions. Multiple representations can be used for content understanding: text descriptions provide a summary of the image or video’s content; audio/visual encoders extract multimodal features from the media; and semantic IDs capture high-level semantic categories. The simulator may also leverage metadata tags, such as hashtags, captions, or timestamps, to better mimic realistic user behavior.
4 Instruction-Following RecSys

### v2

# RecoWorld: Building Simulated Environments for Agentic Recommender Systems

@@ Abstract.
1. Introduction
[…]
| Evaluating RecSys’ Instruction Following Ability. Instruction-following capabilities are essential for agentic RecSys, much like they are for LLMs. E.g., a user might tell the recommender system their preferences through natural language or voice, e.g., “There are too many recommendations about hairstyling; I wanna see something different but related,” or “I’d like to watch more content that people in the San Francisco Bay Area are watching right now.” If the agentic RecSys adjusts its recommendations based on these instructions, we need a simulated environment to evaluate how well it performed. One approach is to measure the system’s success in instruction-following by using simulated user feedback as a proxy. This approach is important because agentic recommenders will need to handle a wide range of user instructions. Our RecoWorld provides a promising way to estimate system performance without relying on ground-truth annotated data. | |
| Enabling Creators to Experiment with Publishing Strategies. A simulated environment helps creators explore different content strategies and reach their goals, e.g., growing followers, boosting views, or getting more likes. It lets them try bold, different approaches without risking real user experience. For example, they can test new topics and styles to attract fresh audiences, find the right posting frequency to keep followers interested without annoying them, or see how controversial opinions might be received. By mimicking user feedback, this environment lets creators test the waters safely. Simulated user profiles can be based on their current audience, a general population sample, or even specific target groups. | |
| Supporting New and Marginal Users in Exploring Interests. Algorithms such as the multi-armed bandit are frequently used in RecSys to find the right balance between exploring new user interests and exploiting known preferences. I.e., an agent chooses from a fixed set of actions (interests) to maximize cumulative reward over time. A challenge here is to explore without compromising user experience. RecoWorld addresses this by collecting simulated user engagement metrics, such as watch time and clicks, which serve as “pseudo-rewards” to evaluate the relevance of each interest. Our simulator can seamlessly integrate with contextual bandit by simulating groups of users with similar features, such as age, location, and past engagements, to provide collective exploration feedback. | |
| Building a Leaderboard for Agentic RecSys. Many enterprise recommender systems are developed in-house, making it difficult to replicate their results. The challenge is even greater with Agentic RecSys, which can take on various forms. They share an “agentic” nature, e.g., an agent (or multiple agents) interacts with an environment; the agent takes actions and receives observations and rewards in return. However, each task may have a different setup, e.g., variations in environmental states and reward signals. RecoWorld provides a shared platform with a community leaderboard, allowing different systems to be compared fairly. We invite practitioners to contribute their agentic tasks to RecoWorld using a common API. They can receive feedback signals from our simulated users and make their tasks available for the community to participate in. RecoWorld thus accelerates RL development, organizes various recommendation tasks from simple to complex for curriculum learning, and even fosters an agent-agent collaborative ecosystem. | |
@@ 1. Introduction
We introduce RecoWorld, a blueprint for building simulated environments that propel the development of agentic recommender systems (Huang et al., 2025a). As these systems gain capabilities such as perception, reasoning, planning, memory, tool use, and autonomy, having environments for rapid system testing and refinement becomes essential before deploying them widely. We do not propose new agentic recommenders here; instead, our focus is on creating simulated environments that replicate user feedback. RecoWorld harnesses LLMs’ strong reasoning, long-context understanding, and multimodal abilities to simulate user responses. It focuses on optimizing for long-term value through simulated interactions. The approach also emphasizes collective impacts, e.g., a post is distributed to a community of interconnected users rather than isolated individuals. The design of user simulation and reward signals in RecoWorld offers promising potential for enhancing existing recommender systems in modeling both short- and long-term user interests, as well as for generative recommenders. (Xia et al., 2025; Wang et al., 2025c).
@@
Our user simulator models sequential user behavior. When a user is about to disengage, they can issue instructions, such as “show me more interesting content”, instead of simply dropping out. The agentic RecSys responds to these instructions by creating new item lists, aiming to re-engage the user. This cycle continues until the user chooses to exit without further instructions.
A successful simulator triggers accurate instructions that guide the recommender, while an effective recommender responds with item lists that enhance user satisfaction. To our knowledge, RecoWorld is the first environment to enable instruction-following recommenders in a Gym-like RL framework, marking a significant step toward more interactive recommender systems.
2. Agentic Environments for RecSys
Agentic is emerging as a defining characteristic of next-generation foundation models. An agentic recommender system acts as an autonomous agent that actively learns from user interactions. It follows user instructions, acquires new skills to make proactive recommendations, and adapts its behavior based on these experiences (Silver and Sutton, 2025). This marks a shift from traditional recommender systems that primarily make passive suggestions (Huang et al., 2025a; Shang et al., 2025; Maragheh and Deldjoo, 2025). With new capabilities to reason and interact with users, there comes the urgent need to develop new environments for evaluation. Simulated environments enable rapid system iteration, enhance data curation, and are increasingly leveraged for agent training. E.g., Claude uses simulated environments to check how the model responds to risky prompts, such as being asked to act as a dark web shopping assistant (Anthropic, 2025).
@@ 2. Agentic Environments for RecSys
Multi-Turn Interactions in a Session. A simulated user interacts with an agentic RecSys over multiple turns (see Figure 1). The user receives a list of recommended items and can take actions such as clicking or skipping. In the first turn, the user clicks on two items (green), skips two items (yellow), and then chooses to leave the session after the next item (red). Before leaving, the user reflects on their experience, identifies reasons for dissatisfaction, and gives a short instruction to the RecSys as feedback. If the RecSys cannot process user instructions, it simply updates recommendations based on recent interactions. If it can, it refreshes the recommendations to better address the user’s feedback. The user reviews the new list, continues interacting, and may repeat the process of providing feedback if considering leaving. This cycle continues until the user decides to end the session without further feedback. The sequence of recommended items and the user’s actions form the interaction trajectory for the session.
Consider an example instruction: User: “Show me more interesting content.” System: “Absolutely! Here are videos tailored to your interests and engagement history. Would you like to specify what you find interesting, such as trending topics, educational content, or entertainment? You can also tell me if you want more from specific creators.” The agent perceives the user’s explicit instruction and reasons over their engagement history, demographics, and behavior to create a configuration for the recommender system to execute. The system returns an updated recommendation list along with strategies to guide the user. The user simulator then generates feedback signals indicating whether the Agentic RecSys met the user’s instructions and may provide the next instruction.
@@
Supporting Real and Simulated User Interactions. (a) When simulated users are replaced by human annotators, the agentic RecSys interacts with actual users and receives authentic feedback signals, effectively functioning as an online experiment. (b) When Agentic RecSys is replaced with a traditional RecSys, the system ignores user instructions and generates recommendations only from the user’s past engagement history, without adapting to additional feedback. (c) When instructions are disabled, the user simulator defaults to an evaluator that scores a list of items, without giving instructions to the RecSys for improvement. Since no improvement is requested, there will be no follow-up iterations; the session will end after a single turn.
Explicit Versus Implicit Instructions Importantly, user preferences are not conveyed solely through explicit instructions; they are also conveyed implicitly through behavior, whether users realize it or not. E.g., if a user frequently finishes watching videos that are ¿5 minutes, it indicates a receptiveness to long-form content, unlike those who prefer quick 10- or 20-second clips. If a user continues a session without interruption, it likely means they are satisfied with the recommendations. Conversely, if they request “something more interesting,” or if a user does not return for a while, it likely indicates that the previous recommendations failed to engage them. Implicit instructions can be inferred by monitoring user state. Users with similar behavioral patterns often share underlying implicit signals. Collaborative filtering can exploit these shared patterns to identify implicit instructions that reflect the collective interests of these user groups. Both implicit and explicit signals are crucial for agentic recommenders.
@@ 3. User Simulation
A session begins when the user opens the app and ends when they exit. Each recommendation presents a list of items, selected from a candidate set and shown in order. The agent and simulated user may interact over multiple turns within a session, e.g., the user requests more interesting content, and the agent updates the list. A reward signal is generated after each list is shown. Crucially, we aim to optimize for long-term user retention as the reward signal, i.e., maximizing session duration and minimizing the interval between sessions, which relates to Daily Active Users (DAU). Specifically, the simulated user goes through each recommendation one by one. For each item, the user takes three steps: 1. Think it through: What is your reasoning behind how you’ll respond? 2. Take action: Write down what you’ll actually do. 3. Update your mindset: How does this affect your current thinking? Table 2 shows a summary containing initial simulator results, including simulated user actions and reasoning for each item.
3.1. Multimodal User Engagment Modeling
@@
Text-based modeling. One approach is to represent all user and item data in textual form . User demographics can be specified in the system prompt, while each item’s metadata, such as publish time, genre, creator, content type, interest cluster, and a brief summary, together with the user’s interaction history, are provided as textual context. Moreover, when short videos are generated directly from user prompts, such as with Sora 2 (OpenAI, 2025), the prompts can be seamlessly incorporated into the item description. This approach offers flexibility in determining which information to include in lifelong user behaviors, supports prompt compression for improved efficiency, and leverages the reasoning ability of LLMs to analyze user behavior and replicate user responses. Nonetheless, relying solely on textual descriptions may overlook certain nuances, including the mood of a video, sarcasm in content, or musical characteristics, which can limit the fidelity of preference modeling.
Multimodal modeling. The second option is to present each item with multimodal information , using a multimodal LLM (MLLM) as the backbone, such as Qwen3’s Omni model (Xu et al., 2025b) or vision-language models (VLM), Gemini-2.5-Pro. In this setting, items are directly ingested by the model rather than represented through textual descriptions. Images are tokenized into patches, while short-form videos with their audio tracks can be jointly processed through frameworks that synchronize visual and acoustic signals. Demographic attributes and textual metadata can also be incorporated as auxiliary inputs. The model output may take multiple modalities, though we restrict it to text for predicting user behavior. Compared with text-only LLMs, multimodal models provide richer representational capacity for multimodal content, enabling more faithful input understanding. Nevertheless, their reasoning ability, typically weaker, and the substantially larger token space of audio–visual inputs imposes significant context and efficiency challenges, making prompt compression particularly difficult.
@@
| Video | Reasoning | Action Taken | Mindset Update |
| 1. Lobster Fishing | Related to fishing interest | Watch (20s) | Open to more fishing content |
| 2. UFC Fight Night | Strong interest, previous engagement | Watch (30s), Like | UFC is a top interest, expect more recommendations |
| 3. Hairstyling | Unrelated, test new content | Skip | Not interested, less likely to engage |
| 4. UFC Fight Night | Interest, but avoid over-engagement | Watch (15s), Skip | Selective with UFC, want variety |
| 5. Hairstyling | Unrelated, already skipped | Skip | Not interested, expect fewer recommendations |
@@
Lifelong User Behavior Modeling. In addition to effective multimodal user engagement modeling, it is essential to achieve robust lifelong user behavior modeling. Users continuously interact with diverse items in recommender systems, resulting in ever-growing interaction histories. Over time, these histories can become effectively unbounded, making it computationally infeasible and costly to store and process all past behaviors. Precisely, we highlight two crucial principles built upon multimodal modeling:
Dynamic memory modeling. Users may have infinite historical interactions . However, not all behaviors contribute equally to reflecting the user’s current intention, and many may introduce noise. It is therefore crucial to design mechanisms that can dynamically and intelligently store, filter, and retrieve user historical behaviors. In RecoWorld, we define an engagement memory at time :
(1) where denotes the action type (e.g., click, like) taken by the user at the -th interaction, is the item interacted with, and is the timestamp of the interaction. is the learned importance score conditioned on current context , and is a retention threshold. This enables dynamic filtering and retrieval of informative past behaviors while discarding noisy or outdated interactions. Session-wise behaviors are also essential to model the user preference. Therefore, we organize the engagement memory at two levels: Interaction-wise memory:
(2) which records fine-grained user actions and the corresponding items . 2) Session-wise memory: We group interactions into sessions , where each session is a trajectory of actions. A session summarization function produces latent representations: and the session-wise memory is defined as:
(3) capturing trajectory-level signals such as mindset shifts and cross-session preference transitions.
Evolving preference modeling. User interests are not static but continuously evolve with changing contexts and the information they are exposed to through recommendations. We denote the user’s latent preference state at time as . Its evolution can be modeled as:
(4) where is the engagement memory (interaction-wise or session-wise), is the current context, and is a temporal update function (e.g., recurrent, attention-based, or diffusion-style). To capture fine-grained preference shifts, we further decompose the update into a mindset update:
(5) where is the mindset adjustment derived from the current memory , context , reasoning trace , and any correction applied by the system to the user’s preference state . This formulation ensures that the simulator captures temporal dynamics, long-term dependencies, and potential preference shifts, such that the generated behaviors remain personalized and timely throughout the interaction lifecycle (see Figure 2).
3.2. Candidate Items
@@
We use real-world content whenever possible (e.g., engagement history and recommended items) to ensure realistic user behavior modeling. Using both simulated content and simulated users can be challenging, as combining these approximations may affect the reliability of behavioral insights. Specifically, each item is presented by its content (e.g., short-form video) along with its metadata and text description. The metadata includes details such as the timestamp, creator information, engagement metrics (e.g., likes, shares, comments), and content category. A user may be recommended various types of content, such as: (a) short-form videos, (b) text posts, (c) mixed media (posts, videos, and images), and (d) friend suggestions. Multiple representations can be used for content understanding: text descriptions provide a summary of the image or video’s content; audio/visual encoders extract multimodal features from the media; and semantic IDs capture high-level semantic categories. The simulator may also leverage metadata tags, such as hashtags or timestamps, to better mimic realistic user behavior.
4. Instruction-Following RecSys